OpenStudy (anonymous):

f(x)= x/(x-5) find the derivative.

OpenStudy (mimi_x3):

\[\frac{d}{dx}\frac{x}{x-5}\] looks like you can use the quotient rule.. or write it as:\[\frac{x}{(x-5)} => x*\frac{1}{(x-5)} => x*(x-5)^{-1}\] To use the product rule

OpenStudy (anonymous):

use the formula (vdu-udv)/v^2

OpenStudy (lgbasallote):

@Mimi_x3 is so smart *_*

OpenStudy (mimi_x3):

lol smart? its piss easy :P ; you're smarter @lgbasallote

OpenStudy (lgbasallote):

hmph dont patronize me @Mimi_x3 you know you're smarter :p